The scalar components of the displacement vector are x = -3 m and y = 4 m. Determine the angle that this vector coincides with the positive direction of the x-axis.


1
Expert's answer
2021-06-07T14:13:18-0400
"\tan\theta=\frac{y}{x}"

"\theta=\tan^{-1}\frac{y}{x}=\tan^{-1}\frac{4}{-3}=127^{\circ}"